A battery with an internal resistance of r and an emf of 10.00 V is connected to a load resistor R=r. As the battery ages, the internal resistance triples. How much is the current through the load resistor reduced?


1
Expert's answer
2021-02-17T10:40:59-0500

Answer

Internal resistance

"i=\frac{E-V}{r+R}"

In first case

"i=\frac{10-V}{r+r}=\frac{10-V}{2r}"

Now in second case

"i'=\frac{10-V}{3r+r}=\frac{10-V}{4r}=\frac{i}{2}"

So current is halves.